There are two small valves in the bottom of a tub filled with water. If opened, one valve will empty the tub in 3 hours; the other will empty it in 6 hours. If both valves are opened at the same time, how long will it take to empty the tub? Please show work.


I have such a problem with algebra. I will appreciate any help you can give me.


Thank you,
Lynnette

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

There are two small valves in the bottom of a tub filled with water. If opened, one valve will empty the tub in 3 hours; the other will empty it in 6 hours. If both valves are opened at the same time, how long will it take to empty the tub? Please show work.
.
Let t = time (hours) to empty tub by both valves
then
t(1/3 + 1/6) = 1
multiplying both sides by 6, we get:
t(2 + 1) = 6
t(3) = 6
t = 6/3
t = 2 hours

Answer by Alan3354(69443)

There are two small valves in the bottom of a tub filled with water. If opened, one valve will empty the tub in 3 hours; the other will empty it in 6 hours. If both valves are opened at the same time, how long will it take to empty the tub? Please show work.
---------------
One drains 1/3 tank/hr
The other drains 1/6 tank/hr
Together, they drain 1/3 + 1/6 tank/hr = 1/2 tank/hour
--> 2 hours per tank